OpenStudy (anonymous):

Suppose you make a solution that contains 13.34 g of sodium chloride and 567g of water. a) What is the mass % of sodium chloride? b) If 11.20mL of your solution weighs 11.364g, what is the density (in g/cc)?

OpenStudy (anonymous):

\[mass~\%~sodium~chloride=\frac{ mass~sodium~chloride }{ mass~solution }=\frac{ 13.34~g }{ 13.34~g+567~g }\] 1 mL = 1 cc, so 11.20 mL = 11.20 cc \[density=\frac{ mass }{ volume }=\frac{ 11.364~g }{ 11.20~cc }\]